The Acme Candy Company claims that 60% of the jawbreakers it produces weigh more than 0.4
ounces. Suppose that 800 jawbreakers are selected at random from the production lines. Use the
Range Rule of Thumb to determine whether it is statistically significant for this sample of 800 to
contain 494 jawbreakers that weigh more than 0.4 ounces? Consider as significant any result that
differs from the mean by more than 2 standard deviations. Show all work (using Range Rule of
Thumb) required to obtain your answer. Clearly state all important values involved.

Answer by Boreal(15235)

The point estimate is 494/800=0.6175
two tail test
z=(phat-p)/sqrt(p*(1-p)/n)
=0.0175/sqrt (.24/800); sqrt term=0.0173
z=1.01
fail to reject Ho because |z|<1.96. The true percentage could be 60%.
p-value=0.31